Find the amount of heat involved when 208.1 g of propane is burned with an excess of oxygen.
C 3H 8(g) + 3 O 2(g) ↔ 3 CO 2(g) + 4 H 2O(l) -->ΔH = -2219 kJ

Key Concept:: Enthalpy is an extensive property, i.e. the magnitude is dependent upon the amount of material involved in the change.
1
Expert's answer
2020-12-18T08:35:37-0500

According to the reaction

"C_3H_{8(g)} + 3 O_{2(g)} ↔ 3CO_{2(g)} + 4 H_2O_{(l)} \qquad ΔH = -2219 kJ/mol"


208.1g of propane = 208.1/44.1 = 4.72 moles


1 mole of propane = 2219kJ evolved


Therefore, 4.72 moles of propane = 2219(4.72) kJ = 10,473 kJ is evolved.